Colin is 9 years older than Richard. David is 12 years younger than Richard. If the total of their ages is 93, how old is the eldest of them?
Answer by ikleyn(52810)

.
Obviously, Colin is the eldest.


Let C be the Colin's age.


Then we have an equation


(C-9) + C + ((C-9)-12) = 93


3C - 30 = 93

3C = 93 + 30 = 123

C = 123/3 = 41.


ANSWER.  The eldest, Colin, is 41 years old.


CHECK.  Then Richard is 41-9 = 32 years old;  David is  32 - 12 = 20 years old  and

        41 + 32 + 20 = 93.   ! Correct !

Solved.